OpenStudy (anonymous):

when solving x^2+4x=-9 what interger or fraction needs to be added to both sides to complete the square

OpenStudy (mertsj):

What is 1/2 the coefficient of the x term?

OpenStudy (anonymous):

Using the base y=ax^2+bx+c, In a perfect square trinomial, (b/2)^2 is c

OpenStudy (anonymous):

I have done (-6)^2-6(1)(-1)/2(1) but after that I am stuck

OpenStudy (anonymous):

x^2+4x=-9 x^2+4x + 4 = -5 (x+2)^2 = -5 I'm not sure what you want to find, so I stop here.